# Material Properties of Popular Radiation Detection Scintillator Crystals for Optical Physics Transport Modelling in Geant4

### v2

@@ I Introduction
Radiation detection is important in a wide range of applications such as environmental monitoring, medical imaging, homeland security, and space [1, 2]. Scintillator-based radiation detectors offer a low power, lightweight, and accurate means to measure radiation. Geant4 [3, 4, 5], a state-of-the-art Monte Carlo radiation transport toolkit that simulates the underlying physical process of interaction and production of particles through matter, has been employed to assist with the design of scintillation detectors [6, 7, 8]. An important aspect of development is accurate scintillator crystal modelling through optimised material properties within Geant4 optical physics. In this work, the material parameters for several common scintillator crystals were used to validate a custom Geant4 (version 11.1.2) experimentally benchmarked silicon photomultiplier (SiPM) based scintillation detector simulation platform. The key material parameters identified to accurately simulate the scintillator crystals were: (1) density, (2) elemental composition, (3) refractive index, (4) optical yield (optical photons produced per MeV of energy deposited), (5) emission spectrum, (6) absorption length, (7) optical decay times, (8) resolution scale, and (9) dimensions [3].
@@ II-A Experimental Platform and Acquisition of Radiation Energy Spectra
The system was placed in a custom-made detector holder (see Fig. (1a)), which was centered on a table away from the lab walls to minimise the chance of Compton scattering with the surrounding environment. The five disk radioactive check sources (109Cd, 57Co, 137Cs, 22Na, and 152Eu) were placed on top of the scintillator crystal within the detector housing to further minimise scattering. After allowing an hour for the SiPM-3000 to thermalise, the detector was calibrated for each crystal type. The integration time, dead time, electronic gain, pulse trigger, and noise trigger were optimised to minimise the noise and low-level detection threshold. For this work, the low-level detection threshold was defined as the minimum deposited gamma ray energy where the corresponding optical photon count exceeded 5% of the primary photopeak count to prevent detector noise from defining the threshold. An integration time of 2 $\mu$s was set for GAGG:Ce, BGO, and NaI:Tl, whereas CLLBC:Ce and CsI:Tl used 3 $\mu$s due to their longer optical decay times (see Table (I)). The SiPM-3000 operating voltage was fixed at 33 V (default), yielding an overvoltage of 4.5 V for the SiPM array. For each scintillator crystal and radioactive source combination, data acquisition ran over 30 minutes in real time and at room temperature (20C) to reduce statistical fluctuations in the energy spectrum and ensure each photopeak could be resolved. Background measurements were also recorded at room temperature over 30 minutes in real time then subtracted from each energy spectrum during post-processing.
@@ II-C Geant4 Application Physics and Optical Surface Modelling
The Geant4 Option4 EM physics list (G4EmStandardPhysics_option4) was used to model the gamma ray, X-ray, and electron transport in the simulation platform [5, 31]. The settings for radioactive decay, atomic de-excitation, PIXE, auger electrons, and fluorescence were enabled. Moreover, the particle production length cut was set to 100 $\mu$m as it was the smallest dimension in the simulation platform. Finally, the low energy cut off was fixed at 100 eV to match the low energy limit of the Geant4 physics model [32]. Optical photon absorption, reflection, and refraction at optical boundaries were modelled using the Geant4 implementation of the Unified model [31, 33]. Since Teflon tape, ESR, and GORE diffuse reflector were highly reflective (see Fig. (8d) in the appendix), optical surfaces in contact with these materials were defined as ‘dielectric-to-metal’. This classification reflects that such surfaces can only absorb and reflect optical photons [34]. In contrast, the remaining surfaces were defined using the ‘dielectric-to-dielectric’ surface type as they can absorb, reflect, and refract optical photons. All optical surfaces were characterised by a ground finish featuring a 0.1-degree surface roughness as it is not possible to have a completely polished surface [15, 34, 6].
@@ III Results and Discussion
The average difference between simulated and experimental photopeak FWHM was $<$1% for CsI:Tl, 1% for GAGG:Ce and CLLBC:Ce, 2% for NaI:Tl, and 4% for BGO. For each scintillator, the energy resolution generally improved as the characteristic gamma ray energy increased. Although this was true for 152Eu within the photopeaks for the same source, it was not true between photopeaks of other isotopes. For example, the 344 keV and 964 keV 152Eu photopeaks for CsI:Tl had a FWHM of 9.55% and 7.17% respectively, however, the FWHM for the 662 keV photopeak was 6.93%. This phenomenon was attributed to cascade summing from multiple 152Eu radioactive decays. The poor energy resolution of BGO blurred the 122 keV and 136 keV 57Co photopeaks together so that the 136 keV photopeak could not be resolved, as shown in Fig. (4b). Moreover, the poor energy resolution of BGO was attributed to the 16% difference between the simulated and experimental 245 keV 152Eu photopeak FWHM. Overall, CLLBC:Ce had the best energy resolution (lowest photopeak FWHM) across the 30 keV to 2 MeV energy range (see Fig. (6)). For the simulated 662 keV photopeak, CLLBC:Ce had an energy resolution of 3.52%, NaI:Tl with 6.70%, GAGG:Ce with 6.83%, CsI:Tl with 6.93%, and BGO with 13.4%.

### v3

@@ I Introduction
Radiation detection is important in a wide range of applications such as environmental monitoring, medical imaging, homeland security, and space [1, 2]. Scintillator-based radiation detectors offer a low power, lightweight, and accurate means to measure radiation. Geant4 [3, 4, 5], a state-of-the-art Monte Carlo radiation transport toolkit that simulates the underlying physical process of interaction and production of particles through matter, has been employed to assist with the design of scintillation detectors [6, 7, 8]. An important aspect of development is accurate scintillator crystal modelling through optimised material properties within Geant4 optical physics. In this work, the material parameters for several common scintillator crystals were used as inputs for a Geant4 (v. 11.1.2) simulation platform designed to model scintillator crystals coupled to a silicon photomultiplier (SiPM) based radiation detector. The key material parameters identified to accurately simulate the scintillator crystals were: (1) density, (2) elemental composition, (3) refractive index, (4) optical yield (optical photons produced per MeV of energy deposited), (5) emission spectrum, (6) absorption length, (7) optical decay times, (8) resolution scale, and (9) dimensions [3].
@@ II-A Experimental Platform and Acquisition of Radiation Energy Spectra
The system was placed in a custom-made detector holder (see Fig. (1a)), which was centered on a table away from the lab walls to minimise the chance of Compton scattering with the surrounding environment. One of each radioactive source (109Cd, 57Co, 137Cs, 22Na, and 152Eu) was placed on top of the scintillator crystal within the detector housing to further minimise scattering. After allowing an hour for the SiPM-3000 to thermalise, the detector was calibrated for each crystal type. The integration time, dead time, electronic gain, pulse trigger, and noise trigger were optimised to minimise the noise and low-level detection threshold. For this work, the low-level detection threshold was defined as the minimum deposited gamma ray energy where the corresponding optical photon count exceeded 5% of the primary photopeak count to minimise the impact of detector noise on the threshold. An integration time of 2 $\mu$s was set for GAGG:Ce, BGO, and NaI:Tl, whereas CLLBC:Ce and CsI:Tl used 3 $\mu$s due to their longer optical decay times (see Table (I)). The SiPM-3000 operating voltage was fixed at 33 V (default), yielding an overvoltage of 4.5 V for the SiPM array. For each scintillator crystal and radioactive source combination, data acquisition ran over 30 minutes in real time and at room temperature (20C) to reduce statistical fluctuations in the energy spectrum and ensure each photopeak could be resolved. Background measurements were also recorded at room temperature over 30 minutes in real time then subtracted from each energy spectrum during post-processing.
@@ II-C Geant4 Application Physics and Optical Surface Modelling
The Geant4 Option4 EM physics list (G4EmStandardPhysics_option4) was used to model the gamma ray, X-ray, and electron transport in the simulation platform [5, 31]. The settings for radioactive decay, atomic de-excitation, PIXE, auger electrons, and fluorescence were enabled. Moreover, the particle production length cut was set to 100 $\mu$m as it was the smallest dimension in the simulation platform. Finally, the low energy cut off was fixed at 100 eV to match the low energy limit of the Geant4 physics model [32]. Optical photon absorption, reflection, and refraction at optical boundaries were modelled using the Geant4 implementation of the Unified model [31, 33]. Optical surfaces were defined in three ways: (1) ‘dielectric-to-dielectric’ with a ground-back-painted finish for surfaces between Teflon and GORE to CLLBC:Ce, NaI:Tl, and CsI:Tl to account for total internal reflection from the air gap, (2) ‘dielectric-to-metal’ with a ground finish for surfaces between ESR and the EPO-TEK-301 glue that binds it to the GAGG:Ce and BGO crystals, and (3) ‘dielectric-to-dielectric’ with a ground finish for the remaining non-reflective material surfaces. All optical surfaces had a 0.1-degree surface roughness as it is not possible to have a completely polished surface [15, 34, 6].
@@ III Results and Discussion
The average difference between simulated and experimental photopeak FWHM was $<$1% for CsI:Tl, 1% for GAGG:Ce and CLLBC:Ce, 2% for NaI:Tl, and 4% for BGO. For each scintillator, the energy resolution generally improved as the characteristic gamma ray energy increased. Although this was true for 152Eu within the photopeaks for the same source, it was not true between photopeaks of other isotopes. For example, the 344 keV and 964 keV 152Eu photopeaks for CsI:Tl had a FWHM of 9.72% and 7.92% respectively, however, the FWHM for the 662 keV photopeak was 7.47%. This phenomenon was attributed to cascade summing from multiple 152Eu radioactive decays. The poor energy resolution of BGO blurred the 122 keV and 136 keV 57Co photopeaks together so that the 136 keV photopeak could not be resolved, as shown in Fig. (4b). Moreover, the poor energy resolution of BGO was attributed to the 16% difference between the simulated and experimental 245 keV 152Eu photopeak FWHM. Overall, CLLBC:Ce had the best energy resolution (lowest photopeak FWHM) across the 30 keV to 2 MeV energy range (see Fig. (6)). For the simulated 662 keV photopeak, CLLBC:Ce had an energy resolution of 3.51%, GAGG:Ce with 6.83%, NaI:Tl with 7.07%, CsI:Tl with 7.47%, and BGO with 13.4%.